OpenStudy (anonymous):

quistion it says The distance traveled is given by the formula d=rt where R is the rate of speed and T is the time spent travling. how long would it take to ride 9 miles at the rate of 6 miles per hour?

OpenStudy (anonymous):

well substitute :) what's the problem?

OpenStudy (anonymous):

its a multiple choice quistion answers are, A) 15 min B) 1.5 hours C) 54 hours D) 54 miles

OpenStudy (anonymous):

yup. so?

OpenStudy (anonymous):

try doing it!

OpenStudy (anonymous):

im not sure how to do it

OpenStudy (anonymous):

is it 54 hours or miles?

OpenStudy (anonymous):

9 times 6 eaquals 54 but im not sure if its miles or hours

OpenStudy (anonymous):

54 hours right?

OpenStudy (anonymous):

so you use dimensions! d = r t . (9 miles) = r * (6 miles / hour) . you can multiply by hours and miles just like any var ;)

OpenStudy (anonymous):

i multiplied them and got 54 is 54 right?

OpenStudy (anonymous):

units. what are the units?

OpenStudy (anonymous):

(9 miles) = (6 miles / hour) *t sorry (error before). uhm. no why would you multiply them?

OpenStudy (anonymous):

sorry im not very good at math im not sure how to do this do i divide?

OpenStudy (anonymous):

....... or multiply by the reciprocal

OpenStudy (anonymous):

so i take 6 times one of my answers then divide by 9

OpenStudy (anonymous):

i think its 1.5

OpenStudy (anonymous):

right?

OpenStudy (phi):

distance traveled is given by the formula D=R*T where R is the rate of speed and T is the time what that sentence says (read it carefully) is DISTANCE is RATE times TIME Now to answer the question how long would it take to ride 9 miles at the rate of 6 miles per hour? how long means TIME 9 miles is DISTANCE 6 miles per hour is RATE in DISTANCE = RATE * TIME replace the things you know: they told you DISTANCE is 9 miles. replace DISTANCE with 9 miles: 9 miles = RATE * TIME they also told you RATE is 6 miles per hour. Replace RATE with 6 miles per hour 9 miles = 6 miles per hour * TIME now you need to find TIME. The idea is divide both sides by 6 miles per hour \[ \frac{ 9 \text{ miles}}{6 \text{ miles/hour}} = \frac{6 \text{ miles/hour}*TIME}{6 \text{ miles/hour}} \] the 6 mph on the right side divide out (anything divided by itself is 1)

OpenStudy (anonymous):

k thxs